# A Perfectly Truthful Calibration Measure

### v1

@@ 1.3 Our Contributions
Our construction of $\mathrm{ATB}$ is a special case of UBSE, and thus the truthfulness of $\mathrm{ATB}$ (Theorem 1.3) and the rank-preserving property (Theorem 1.6) follow immediately from Lemma 1.7. We remark that while every UBSE is truthful, it may not have the other desired properties of ATB, such as soundness, continuity, and the relationship to smCal and $\underline{\textsc{distCal}}$.
@@
Faster Optimal Calibration Test via Validity.
@@
Given $T$ samples, $\mathrm{ATB}$ can distinguish a calibrated predictor from a predictor with $\textsc{smCal}\geq C/\sqrt{T}$ for an absolute constant $C>0$. Moreover, this rate is information-theoretically optimal (up to constant factors). This also holds when we replace smCal with $\underline{\textsc{distCal}}$ because the two are constant-factor approximations of each other as shown by Błasiok et al. (2023).
As mentioned earlier, ATB can be easily computed in $O(T\log T)$ time. Thus, Theorem 1.8 implies a faster calibration tester than the previous $O(T\log^{2}T)$-time calibration tester for smCal and $\underline{\textsc{distCal}}$ from Hu et al. (2024).
@@
We establish the quadratic relationship between $\mathrm{ATB}$ and $\textsc{smCal},\underline{\textsc{distCal}}$ using the following $\ell_{1}$ variant of $\mathrm{ATB}$. We show that this variant gives a constant-factor approximation for smCal and $\underline{\textsc{distCal}}$.
@@
Similar to $\mathrm{ATB}$, $\ell_{1}\textup{-}\mathrm{ATB}$ can also be easily computed in $O(T\log T)$ time (see Section 4.4), and thus gives a faster algorithm for approximating smCal and $\underline{\textsc{distCal}}$ than direct computation, for which the best-known algorithms take $O(T\log^{2}T)$ and $O(T^{2}\log T)$ time, respectively (Hu et al., 2024).
@@ 1.4.1 Truthfulness from Variance Additivity
However, $\ell_{2}\textup{-}\textsc{BinECE}$ is not truthful because it does not bin based on indices $t$, but rather the prediction values $r_{t}$. Specifically, $\ell_{2}\textup{-}\textsc{BinECE}$ combines adjacent predictions into the same bin. It works by first partitioning the prediction space $[0,1]$ into intervals, with one bin corresponding to each interval. Each bin is decided by the predictions $v_{t}$ that fall within the corresponding interval. A strategic predictor will try to reduce the expected error by manipulating the partition $\mathcal{B}$ via making untruthful predictions. For the same Example 1.1, always predicting the same value $r_{1}=\cdots=r_{T}$ puts all the indices in the same bin, resulting in a different index partition $\mathcal{B}^{\prime}$ than predicting truthfully. This different partition can significantly reduce the expected $\ell_{2}\textup{-}\textsc{BinECE}$ due to the bin-size-based normalization $1/|B_{i}|$ in (2).
The example above hints that we can achieve truthfulness if the expected error of truthful predictions is invaiant to the index partition $\mathcal{B}$. To see this, our analysis above shows that for any partition $\mathcal{B}^{\prime}$ that could potentially be manipulated from a strategic report $\bm{r}$, predicting the truth achieves smaller or equal expected error on the same parition $\mathcal{B}^{\prime}$:
@@
The final observation that completes our construction of UBSEs is that, by the linearity of expectation, we can even allow the partition $\mathcal{B}$ to be randomized and use the expected error over the random partition $\mathcal{B}$, without breaking the truthfulness analysis above. This allows us to construct a general family of truthful calibration errors, each using a different, possibility randomized, partition $\mathcal{B}$ that can depend on the predictions $r_{1},\ldots,r_{T}$. We term these calibration errors Unnormalized Binned Squared Errors (UBSEs) and present the formal definition and analysis in Section 3.
@@
  • Divide predictions into $k$ bins, with $\frac{T}{k}$ numbers of samples in each bin.
@@ 1.4.2 Two-Bin Approximation of the Smooth Calibration Error
While neither inequality is straightforward to prove, the relatively more technically involved and, perhaps, more surprising direction is the latter inequality (9) showing that smCal can be upper-bounded by $\ell_{1}\textup{-}\mathrm{ATB}$ up to a constant factor. Indeed, the intuition behind the previous notion of interval calibration error intCal (Błasiok et al., 2023) is that having too few bins tends to underestimate smCal, and if the calibration error is much smaller than the average bin width, we should increase the number of bins to faithfully capture smCal. [^3] Consequently, the number of bins used to define $\textsc{intCal}(\bm{r},\bm{y})$ depends on both $\bm{r}$ and $\bm{y}$. In UBSE, the binning scheme can only depend on $\bm{r}$ in order for our truthfulness analysis to hold. The reasoning is that having fewer bins makes more predictions fall into the same bin, among which the biases $r_{t}-y_{t}$ in positive and negative directions cancel out, thus more likely to cause underestimation. For example, having only one bin gives the following UBSE:
@@
To fully prove (9), we need to remove the monotonicity and differentiability assumptions on $w$, which is achieved by our formal proof in Section 5. Roughly speaking, without monotonicity, the convex combination of the threshold functions that expresses $w$ might have negative coefficients (so it is a linear combination rather than a convex combination), but the absolute values of the coefficients can still be controlled using the Lipschitzness of $w$. The differentiability assumption can be removed by focusing on the finite set $\{r_{1},\ldots,r_{T}\}$ rather than the full domain $[0,1]$ of $w$.
@@ 1.7 Paper Organization
The rest of the paper is organized as follows. Section 2 establishes the basic setup, including the definitions of existing calibration errors (Section 2.1), completeness and soundness (Section 2.2), the validity of calibration errors via calibration tests (Section 2.3), and the truthfulness of calibration errors (Section 2.4). In Section 3, we introduce the Unnormalized Binned Squared Errors (UBSEs), a general family of truthful binning-based error metrics. In Section 4, we introduce our proposed calibration error, the Averaged Two-Bin calibration error ($\mathrm{ATB}$) as a special case of UBSE and prove its truthfulness, continuity, sample efficiency, and computationally efficiency. In Section 5, we prove the quadratic relationship between $\mathrm{ATB}$ and the existing calibration errors $\textsc{smCal},\underline{\textsc{distCal}}$ by showing that $\ell_{1}\textup{-}\mathrm{ATB}$ is a constant-factor approximation of smCal and $\underline{\textsc{distCal}}$. In Section 6, we show that $\mathrm{ATB}$ is optimally valid for smCal and $\underline{\textsc{distCal}}$, implying a simple and efficient calibration tester for smCal and $\underline{\textsc{distCal}}$ with improved running time over previous work.
@@ 2.2 Completeness and Soundness over Samples
We can estimate $\ell_{\alpha}\textup{-}\textsc{BinECE}(J)$ using a sample $S=\{(v_{t},y_{t})\}_{t\in T}$ of $T$ i.i.d. points drawn from $J$. Specifically, letting $J_{S}$ denote the (empirical) uniform distribution over $S$, we can use $\ell_{\alpha}\textup{-}\textsc{BinECE}(J_{S})$ as a good estimate for $\ell_{\alpha}\textup{-}\textsc{BinECE}(J)$ when the sample size $T$ is sufficiently large relative to $k$ (the number of bins). In practice, the number $k$ of bins can be selected according to the sample size $T$, e.g. $k=T^{\frac{1}{3}}$.
@@
We see that the error $\textsc{Cal}=({\mathbf{E}}\left[r\right]-{\mathbf{E}}\left[y\right])^{2}+{\mathbf{E}}\left[\mathbb{I}\left[r\in\{0,1\},y\neq r\right]\right]$ satisfies the requirements above. However, for predictions not in $\{0,1\}$, the error metric only evaluates the unconditional bias in predictions, which is far from a calibration error metric.
@@ 2.3 Calibration Test and Validity
2.4 Truthfulness
A truthful error metric incentivizes a strategizing predictor to report the true distribution to minimize expected error on a finite sample.
Definition 2.14 (Truthfulness on Samples).
We say a calibration measure Cal is truthful on samples if the following holds. Let $D$ be an arbitrary joint distribution of $(x,y)\in X\times\{0,1\}$ and let $p:X\to[0,1]$ be the ground-truth predictor $p(x)=\operatorname{\mathbb{E}}_{D}[y|x]$. Let $S=\{(x_{t},y_{t})\}_{t\in T}$ be a sample of $T$ i.i.d. points drawn from $D$, and let $D_{S}$ denote the (empirical) uniform distribution over $S$. Then
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)]\leq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)]\quad\text{for any predictor $r:X\to[0,1]$}.$
In this paper, we focus on a stronger notion of truthfulness, where the predictor makes predictions for an arbitrary set of $T$ individuals that may not be i.i.d. sampled from any underlying distribution. We first extend our definition of calibration errors to this setting, where we evaluate the calibration error of a reported prediction sequence $\bm{r}=(r_{1},\ldots,r_{T})$ for the $T$ individuals w.r.t. a ground-truth probability sequence $\bm{p}=(p_{1},\ldots,p_{T})$.
$\textsc{Cal}(\bm{r},\bm{p}):=\textsc{Cal}(J_{\bm{r},\bm{p}}).$
We have
$\textsc{ECE}(\bm{r},\bm{p})=\textsc{ECE}(J_{\bm{r},\bm{p}})$ $={\mathbf{E}}_{(v,y)\sim J_{\bm{r},\bm{p}}}\left[\big{|}v-\widehat{v}\big{|}\right$]
$=\frac{1}{T}\sum_{v}\sum_{t\in[T]}\mathbb{I}\left[r_{t}=v\right]\big{|}v-\widehat{v}\big{|}$ ($v$ ranges over all values that appear at least once in the set $\{r_{1},\ldots r_{t}\}$)
Similarly for smCal:
Definition 2.16 (Truthfulness on Sequences).
We say a calibration measure Cal is truthful on sequences if the following holds for any $T\in\mathbb{Z}_{>0}$. Let $\bm{p}:=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ be an arbitrary sequence of ground-truth predictions. Let $\bm{y}=(y_{1},\ldots,y_{t})$ denote the randomly realized states, where each $y_{t}\in\{0,1\}$ is drawn independently from the Bernoulli distribution with mean $p_{i}$ (denoted $\bm{y}\sim\bm{p}$). Then
Claim 2.17.
Truthfulness on sequences implies truthfulness on samples.
Proof.
Assuming truthfulness on sequences, we know that the quantity in (17) is no smaller than the quantity in (18). Taking the expectation over $x_{1},\ldots,x_{T}$ proves the desired truthfulness on samples. ∎
3 Truthfulness of Unnormalized Binned Squared Errors
@@
This small difference makes UBSEs truthful (whereas the binned $\ell_{2}\text{-}\textsc{ECE}$ is not):
@@
Any UBSE error metric Cal is truthful on sequences (Definition 2.16).
@@
Clearly, we have $\textsc{Cal}(\bm{r},\bm{p})\geq 0$ and $\textsc{Cal}(\bm{p},\bm{p})=0$. Therefore, the quantity in (20) is non-negative, which means that Cal is truthful on sequences. ∎
@@
[…]

### v2

@@ 1.3 A Roadmap of Our Contributions
To see the truthfulness, when $\bm{r}=\bm{p}$, the middle term is 0, while the third term is invariant of the predictions. The expected error is thus minimized at $\bm{r}=\bm{p}$.
Our construction of $\mathrm{ATB}$ is a special case of UBSE, and thus the truthfulness of $\mathrm{ATB}$ follows immediately from Lemma 1.3. We remark that while every UBSE is truthful, it may not have the other desired properties of ATB, such as soundness, continuity, and the relationship to smCal and $\underline{\textsc{distCal}}$.
@@
Implications of Truthfulness.
Section 5 discusses the implications of a truthful error measure beyond the definition of truthfulness. Following the literature on proper (a.k.a., truthful) loss functions (Gneiting, 2011), we characterize the ranking of predictors by the truthful expected error.
  • Theorem 5.2 shows, among calibrated predictors, a truthful error measure consistently ranks a Blackwell dominating predictor with a (weakly) lower expected error. A Blackwell dominating predictor is more informative and helps any decision maker achieve a lower loss regardless of the loss function. For example, the truth-telling predictor dominates all other predictors.
  • Theorem 5.3 shows, for every miscalibrated predictor, calibrating the predictor always (weakly) reduces the expected truthful error.
Results in Section 5 hold generally for every truthful error measure, not just for sound and complete calibration measures.
Linear-Time Calibration Test via Validity.
@@
Given $T$ examples, $\mathrm{ATB}$ can distinguish a calibrated predictor from a predictor with $\textsc{smCal}\geq C/\sqrt{T}$ for an absolute constant $C>0$. Moreover, this rate is information-theoretically optimal (up to constant factors). This also holds when we replace smCal with $\underline{\textsc{distCal}}$ because the two are constant-factor approximations of each other as shown by Błasiok et al. (2023).
As mentioned earlier, ATB can be computed in $O(T\log T)$ time. Thus, Theorem 1.4 implies a faster calibration tester than the previous $O(T\log^{2}T)$-time calibration tester for smCal and $\underline{\textsc{distCal}}$ from Hu et al. (2024). Moreover, the running time can be further improved to $O(T)$ because we do not need to compute ATB exactly and just need to approximate it up to error $O(1/T)$ (see Section 6). This gives the first linear-time algorithm for calibration testing.
@@
As mentioned above, the completeness and soundness are proven by the quadratic relationship of ATB to $\textsc{smCal},\underline{\textsc{distCal}}$. This relationship is established via the following non-truthful $\ell_{1}$ variant of $\mathrm{ATB}$, which quadratically approximates $\mathrm{ATB}$. We show that this variant gives a constant-factor approximation for smCal and $\underline{\textsc{distCal}}$.
@@
Similar to $\mathrm{ATB}$, $\ell_{1}\textup{-}\mathrm{ATB}$ can also be easily computed in $O(T\log T)$ time and approximated up to $\varepsilon>0$ additive error in $O(T+1/\varepsilon)$ time (see Section 4.5). Therefore, Theorem 1.6 implies a faster constant-factor approximation algorithm for smCal and $\underline{\textsc{distCal}}$ than direct computation, for which the best-known algorithms take $O(T\log^{2}T)$ and $O(T^{2}\log T)$ time, respectively (Hu et al., 2024).
@@ 1.4.1 Truthfulness from Variance Additivity
However, $\ell_{2}\textup{-}\textsc{BinECE}$ in its standard form is not truthful because it does not bin based on indices $t$, but rather the prediction values $r_{t}$. Specifically, $\ell_{2}\textup{-}\textsc{BinECE}$ combines adjacent predictions into the same bin. It works by first partitioning the prediction space $[0,1]$ into intervals, with one bin corresponding to each interval. Each bin consists of the predictions $v_{t}$ that fall within the corresponding interval. A strategic predictor will try to reduce the expected error by manipulating the partition $\mathcal{B}$ via making untruthful predictions. For the same Example 1.1, always predicting the same value $r_{1}=\cdots=r_{T}$ puts all the indices in the same bin, resulting in a different index partition $\mathcal{B}^{\prime}$ than predicting truthfully. This different partition can significantly reduce the expected $\ell_{2}\textup{-}\textsc{BinECE}$ due to the bin-size-based normalization $1/|B_{i}|$ in (2).
The example above hints that we can achieve truthfulness if the expected error of truthful predictions is invaiant to the index partition $\mathcal{B}$. To see this, our analysis above shows that for any index partition $\mathcal{B}^{\prime}$ that could potentially be manipulated from a strategic report $\bm{r}$, predicting the truth achieves smaller or equal expected error on the same index partition $\mathcal{B}^{\prime}$:
@@
The final observation that completes our construction of UBSEs is that, by linearity of expectation, we can even allow the partition $\mathcal{B}$ to be randomized and use the expected error over the random partition $\mathcal{B}$, without breaking the truthfulness analysis above. This allows us to construct a general family of truthful calibration errors, each using a different, possibility randomized, partition $\mathcal{B}$ that can depend on the predictions $r_{1},\ldots,r_{T}$. We term these calibration errors Unnormalized Binned Squared Errors (UBSEs) and present the formal definition and analysis in Section 3.
@@
  • Divide predictions into $k$ bins, with $\frac{T}{k}$ predictions in each bin.
@@ 1.4.2 Two-Bin Approximation of the Smooth Calibration Error
While neither inequality is straightforward to prove, the relatively more technically involved and, perhaps, more surprising direction is the latter inequality (9) showing that smCal can be upper-bounded by $\ell_{1}\textup{-}\mathrm{ATB}$ up to a constant factor. Indeed, the intuition behind the previous notion of interval calibration error intCal (Błasiok et al., 2023) is that having too few bins tends to underestimate smCal, and if the calibration error is much smaller than the average bin width, we should increase the number of bins to faithfully capture smCal. [^3] Consequently, the number of bins used to define $\textsc{intCal}(\bm{r},\bm{y})$ depends on both $\bm{r}$ and $\bm{y}$. In UBSE, the binning scheme can only depend on $\bm{r}$ in order for our truthfulness analysis to hold. The reasoning is that having fewer bins makes more predictions fall into the same bin, among which the positive and negative biases $r_{t}-y_{t}$ cancel out, thus more likely to cause underestimation. For example, having only one bin gives the following UBSE:
@@
To fully prove (9), we need to remove the monotonicity and differentiability assumptions on $w$, which is achieved by our formal proof in Section 4.1.1. Roughly speaking, without monotonicity, the convex combination of the threshold functions that expresses $w$ might have negative coefficients (so it is a linear combination rather than a convex combination), but the absolute values of the coefficients can still be controlled using the Lipschitzness of $w$. The differentiability assumption can be removed by focusing on the finite set $\{r_{1},\ldots,r_{T}\}$ rather than the full domain $[0,1]$ of $w$.
@@ 1.7 Paper Organization
The rest of the paper is organized as follows. Section 2 establishes the basic setup, including the definitions of existing calibration errors (Section 2.1), completeness and soundness (Section 2.2), the validity of calibration errors via calibration tests (Section 2.4), and the truthfulness of calibration errors (Section 2.3). In Section 3, we introduce the Unnormalized Binned Squared Errors (UBSEs), a general family of truthful binning-based error metrics. In Section 4, we introduce our proposed calibration error, the Averaged Two-Bin calibration error ($\mathrm{ATB}$) as a special case of UBSE and prove its truthfulness, continuity, sample efficiency, and computationally efficiency. In Section 4.1.1, we prove the quadratic relationship between $\mathrm{ATB}$ and the existing calibration errors $\textsc{smCal},\underline{\textsc{distCal}}$ by showing that $\ell_{1}\textup{-}\mathrm{ATB}$ is a constant-factor approximation of smCal and $\underline{\textsc{distCal}}$. Section 5 discusses the implications of truthfulness: preservation of Blackwell ordering and monotonicity under recalibration, as well as robustness to hyperparameter selection. In Section 6, we show that $\mathrm{ATB}$ is optimally valid for smCal and $\underline{\textsc{distCal}}$, implying a linear-time calibration tester for smCal and $\underline{\textsc{distCal}}$.
@@ 2 Preliminaries
We present useful definitions and preliminary theorems for our paper. Section 2.1 introduces existing calibration measures. Section 2.2 defines the completeness and soundness of a calibration measure. Section 2.3 formalizes truthfulness of an error measure. Section 2.4 introduces calibration test, preparing for the result on linear-time calibration tester.
@@ 2.2 Completeness and Soundness over Samples
We can estimate $\ell_{\alpha}\textup{-}\textsc{BinECE}(J)$ using a sample $S=\{(v_{t},y_{t})\}_{t\in T}$ of $T$ i.i.d. points drawn from $J$. Specifically, letting $J_{S}$ denote the (empirical) uniform distribution over $S$, we can use $\ell_{\alpha}\textup{-}\textsc{BinECE}(J_{S})$ as a good estimate for $\ell_{\alpha}\textup{-}\textsc{BinECE}(J)$ when the sample size $T$ is sufficiently large relative to $k$ (the number of bins). In practice, the number $k$ of bins can be selected according to the sample size $T$, e.g. $k=T^{\frac{1}{3}}$, to ensure soundness and completeness.
@@
We see that the error $\textsc{Cal}=({\operatorname{\mathbb{E}}}\left[r\right]-{\operatorname{\mathbb{E}}}\left[y\right])^{2}+{\operatorname{\mathbb{E}}}\left[\mathbb{I}\left[r\in\{0,1\},y\neq r\right]\right]$ satisfies the requirements above. However, for predictions not in $\{0,1\}$, the error metric only evaluates the unconditional bias in predictions, which is far from a calibration error metric.
2.3 Truthfulness
A truthful error metric incentivizes a strategizing predictor to report the true distribution to minimize expected error on a finite sample. Definition 2.12 defines the ex-ante truthfulness where a predictor output is assumed to be function of the feature space.
Definition 2.12 (Ex-Ante Truthfulness).
We say a calibration measure Cal is ex-ante truthful if the following holds. Let $D$ be an arbitrary joint distribution of $(x,y)\in X\times\{0,1\}$ and let $p:X\to[0,1]$ be the ground-truth predictor $p(x)=\operatorname{\mathbb{E}}_{D}[y|x]$. Let $S=\{(x_{t},y_{t})\}_{t\in[T]}$ be a sample of $T$ i.i.d. points drawn from $D$, and let $D_{S}$ denote the (empirical) uniform distribution over $S$. Then
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)]\leq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)]\quad\text{for every predictor $r:X\to[0,1]$}.$
In this paper, we study a strictly stronger notion: interim truthfulness. In the interim stage, the true distribution of $T$ samples are realized, and the predictor is allowed to deviate and report any prediction sequence, not necessarily a function of the feature space. We first extend our definition of calibration errors to this setting, where we evaluate the calibration error of a reported prediction sequence $\bm{r}=(r_{1},\ldots,r_{T})$ for the $T$ individuals w.r.t. a ground-truth probability sequence $\bm{p}=(p_{1},\ldots,p_{T})$. We will refer to both definition as truthfulness when it is clear from the context.
$\textsc{Cal}(\bm{r},\bm{p}):=\textsc{Cal}(J_{\bm{r},\bm{p}}).$
We have
[…]
Similarly for smCal:
Definition 2.14 (Interim Truthfulness).
We say a calibration measure Cal is interim truthful if the following holds for any $T\in\mathbb{Z}_{>0}$. Let $\bm{p}:=(p_{1},\ldots,p_{T})\in[0,1]^{T}$ be an arbitrary sequence of ground-truth predictions. Let $\bm{y}=(y_{1},\ldots,y_{t})$ denote the randomly realized states, where each $y_{t}\in\{0,1\}$ is drawn independently from the Bernoulli distribution with mean $p_{i}$ (denoted $\bm{y}\sim\bm{p}$). Then
Claim 2.15 (Interim truthfulness implies ex-ante truthfulness).
Let $\textsc{Cal}(\bm{r},\bm{p})$ be a calibration measure induced by $\textsc{Cal}(J)$ (Definition 2.13). If $\textsc{Cal}(\bm{r},\bm{p})$ is interim truthful, then $\textsc{Cal}(J)$ is ex-ante truthful.
Proof.
Assuming interim truthfulness, we know that the quantity in (17) is no smaller than the quantity in (18). Taking the expectation over $x_{1},\ldots,x_{T}$ proves the desired ex-ante truthfulness. ∎
@@ 2.4 Calibration Test and Validity
3 Truthful Family: Unnormalized Binned Squared Errors
@@
3.1 Interim Truthfulness
UBSEs is interim truthful (whereas the binned $\ell_{2}\text{-}\textsc{ECE}$ is not, with the small difference above):
@@ 3.1 Interim Truthfulness
Any UBSE error metric Cal is interim truthful (Definition 2.14).
@@
Clearly, we have $\textsc{Cal}(\bm{r},\bm{p})\geq 0$ and $\textsc{Cal}(\bm{p},\bm{p})=0$. Therefore, the quantity in (20) is non-negative, which means that Cal is interim truthful. ∎
@@
In (21), we used the fact that the distribution of $\mathcal{B}$ depends only on $\bm{r}$ and not on $\bm{y}$. For the same reason, the two distributions of $\mathcal{B}$ in (21) and (22) are the same. Therefore, to prove the lemma, it suffices to show that for any fixed partition $\mathcal{B}$,
@@
3.2 Strict Ex-Ante Truthfulness
In the ex-ante stage before the ground-truth probabilities for each sample are drawn, UBSE is strictly truthful, i.e., the unique minimizer to the expected error is the ground truth predictor $p(x):=\operatorname{\mathbb{E}}[y|x]$.
Theorem 3.5.
Let $X$ be an arbitrary non-empty domain and let $D$ be an arbitrary distribution of $(x,y)\in X\times\{0,1\}$. Let $p:X\to[0,1]$ be the ground-truth predictor $p(x):=\operatorname{\mathbb{E}}[y|x]$ and let $r:X\to[0,1]$ be an arbitrary predictor. Let Cal be an arbitrary UBSE and $T$ be an arbitrary positive integer. For a sample $S=((x_{t},y_{t}))_{t\in[T]}$ of $T$ i.i.d. examples drawn from $D$, let $D_{S}$ denote the uniform distribution on $S$. Suppose
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)]\leq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)].$
Then $r(x)=p(x)$ holds almost surely over $(x,y)\sim D$.
We prove Theorem 3.5 using the following lemma. We are able to prove a stronger version of this lemma for the special case of $\mathrm{ATB}$ in Lemma C.1.
Lemma 3.6.
Let $J$ be an arbitrary distribution of $(v,y)\in[0,1]\times\{0,1\}$, and define random variable $\widehat{v}:=\operatorname{\mathbb{E}}_{J}[y|v]$ as a function of $v$. Let Cal be an arbitrary UBSE. For $T$ i.i.d. examples $(\widehat{v}_{1},v_{1},y_{1}),\ldots,\allowbreak(\widehat{v}_{T},v_{T},y_{T})$, defining $\widehat{\bm{v}}:=(\widehat{v}_{1},\ldots,\widehat{v}_{T}),\bm{v}:=(v_{1},\ldots,v_{T})$ and $\bm{y}=(y_{1},\ldots,y_{T})$, we have
$\operatorname{\mathbb{E}}[\textsc{Cal}(\widehat{\bm{v}},\bm{y})]$ $=\frac{1}{T}\operatorname{\mathbb{E}}_{J}[(\widehat{v}-y)^{2}],$ (25)
$\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{v},\bm{y})]$ $\geq\operatorname{\mathbb{E}}[\textsc{Cal}(\widehat{\bm{v}},\bm{y})]+\frac{1}{T^{2}}\operatorname{\mathbb{E}}_{J}[(\widehat{v}-v)^{2}].$ (26)
Proof.
By the definition of $\widehat{v}:=\operatorname{\mathbb{E}}_{J}[y|v]$, the distribution of $(\widehat{v},y)$ is calibrated. Therefore, $\operatorname{\mathbb{E}}[y_{t}|\widehat{v}_{t}]=\widehat{v}_{t}$ for every $t=1,\ldots,T$. By Lemma 3.3, we have
$\operatorname{\mathbb{E}}[\textsc{Cal}(\widehat{\bm{v}},\bm{y})]=\frac{1}{T^{2}}\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\widehat{v}_{t}(1-\widehat{v}_{t})\right]=\frac{1}{T}\operatorname{\mathbb{E}}[\widehat{v}(1-\widehat{v})]=\frac{1}{T}\operatorname{\mathbb{E}}[(\widehat{v}-y)^{2}],$
where we use the fact that $\widehat{v}_{1},\ldots,\widehat{v}_{T}$ are i.i.d. random variables. This completes the proof of (25).
By Lemma 3.3 again, we have
$\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{v},\bm{y})]=\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{v},\widehat{\bm{v}})]+\frac{1}{T^{2}}\operatorname{\mathbb{E}}\left[\sum_{t=1}^{T}\widehat{v}_{t}(1-\widehat{v}_{t})\right]=\operatorname{\mathbb{E}}[\textsc{Cal}(\bm{v},\widehat{\bm{v}})]+\operatorname{\mathbb{E}}[\textsc{Cal}(\widehat{\bm{v}},\bm{y})].$ (27)
Since Cal is a UBSE (Definition 3.1) and there are at most $T$ non-empty bins, by the Cauchy-Schwarz inequality, we have
$\textsc{Cal}(\bm{v},\widehat{\bm{v}})\geq\frac{1}{T^{3}}\left(\sum_{t=1}^{T}(v_{t}-\widehat{v}_{t})\right)^{2}.$
Taking expectation, we get
[…]
$\geq\frac{1}{T^{2}}\operatorname{\mathbb{E}}[(v-\widehat{v})^{2}].$
Plugging this into (27) proves (26). ∎
We are now ready to prove Theorem 3.5.
Proof of Theorem 3.5.
Define predictor $\widehat{r}:X\to[0,1]$ such that $\widehat{r}(x)=\operatorname{\mathbb{E}}_{D}[y|r(x)]$. By Lemma 3.6,
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)]$ $=\frac{1}{T}\operatorname{\mathbb{E}}_{D}[(p(x)-y)^{2}],$
$\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)]$ $\geq\frac{1}{T}\operatorname{\mathbb{E}}_{D}[(\widehat{r}(x)-y)^{2}]+\frac{1}{T^{2}}\operatorname{\mathbb{E}}_{D}[(r(x)-\widehat{r}(x))^{2}].$
Since $\operatorname{\mathbb{E}}[y|x]=p(x)$, we have
$\operatorname{\mathbb{E}}_{D}[(\widehat{r}(x)-y)^{2}]$ $=\operatorname{\mathbb{E}}[(\widehat{r}(x)-p(x))^{2}]+\operatorname{\mathbb{E}}[(p(x)-y)^{2}]+2\operatorname{\mathbb{E}}[(\widehat{r}(x)-p(x))(p(x)-y)]$
$=\operatorname{\mathbb{E}}[(\widehat{r}(x)-p(x))^{2}]+\operatorname{\mathbb{E}}[(p(x)-y)^{2}].$
Therefore,
$0$ $\geq\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(r)]-\operatorname{\mathbb{E}}_{S}[\textsc{Cal}_{D_{S}}(p)]$
$\geq\frac{1}{T}(\operatorname{\mathbb{E}}[(\widehat{r}(x)-y)^{2}]-\operatorname{\mathbb{E}}[(p(x)-y)^{2}])+\frac{1}{T^{2}}\operatorname{\mathbb{E}}[(r(x)-\widehat{r}(x))^{2}]$
$=\frac{1}{T}\operatorname{\mathbb{E}}[(\widehat{r}(x)-p(x))^{2}]+\frac{1}{T^{2}}\operatorname{\mathbb{E}}[(r(x)-\widehat{r}(x))^{2}].$
This implies that $r(x)=\widehat{r}(x)=p(x)$ almost surely. ∎